Please help me solve this:
The width of a rectangle is 3 less than twice the length x. If the area of the rectangle is 43, which equation can be used to find the length, in feet?
1. 2x (x-3)=43
2. x(3-2x) =43
3. x(2x - 3)=43
4. 2x+2(2x-3)=43
P.S: I know the answer is #3 but I don't know how to prove it with work!

Answer by jim_thompson5910(35256)

Let x be the length

"The width of a rectangle is 3 less than twice the length x", so the length is 2x-3

-------------------------------------------------------

Area = (Length)*(Width)

Area = (2x-3)*(x)

Area = x*(2x-3)

43 = x*(2x-3)

x*(2x-3) = 43